OpenStudy (anonymous):

What is 4(x-2) = 10x-40?

OpenStudy (anonymous):

4x -8 = 10x-40 4x - 10x = -40 + 8 -6x = -32/:-6 x=5 and 2/6

OpenStudy (amistre64):

it would be better if you allow the asker to participate in the solution process

OpenStudy (anonymous):

Thank you!

OpenStudy (radicalsniper):

Combine the like terms :)